OpenStudy (redneckoutlaw):

What is the product of 5the square root of 25 times d. • 2dthe square root of 3.? Simplify if possible.

OpenStudy (anonymous):

\[\large \sqrt[5]{25}\times \sqrt{3}\]??

OpenStudy (anonymous):

or \[5\sqrt{25}\times \sqrt3\]

OpenStudy (redneckoutlaw):

its 2√3

OpenStudy (anonymous):

not clear try the equation tool

OpenStudy (redneckoutlaw):